QUESTION

Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $120 per unit. Variable expenses are $60.00 per unit, and fixed expenses total $160,000 per year. Its operating results for last year were as follows:

Sales $ 2,880,000

Variable expenses 1,440,000

Contribution margin 1,440,000

Fixed expenses 160,000

Net operating income $ 1,280,000

1.) The sales manager is convinced that a 11% reduction in the selling price, combined with a $65,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

2.) The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.00 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $1,280,000 net operating income as last year?

QUESTION

Northwood Company manufactures basketballs. The company has a ball that sells for $25. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $15.00 per ball, of which 60% is direct labor cost. Last year, the company sold 32,000 of these balls, with the following results:

Sales (32,000 balls) $ 800,000

Variable expenses 480,000

Contribution margin 320,000

Fixed expenses 211,000

Net operating income $ 109,000

Required:

1. Compute (a) last year's CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last year’s sales level.

2. Due to an increase in labor rates, the company estimates that next year's variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $25.00, what will be next year's CM ratio and the break-even point in balls?

3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next year to earn the same net operating income, $109,000, as last year?

4. Refer again to the data in (2) above. The president feels that the company must raise the selling price of its basketballs. If Northwood Company wants to maintain the same CM ratio as last year (as computed in requirement 1a), what selling price per ball must it charge next year to cover the increased labor costs?

5. Refer to the original data. The company is discussing the construction of a new, automated manufacturing plant. The new plant would slash variable expenses per ball by 40.00%, but it would cause fixed expenses per year to double. If the new plant is built, what would be the company’s new CM ratio and new break-even point in balls?

6. Refer to the data in (5) above.

a. If the new plant is built, how many balls will have to be sold next year to earn the same net operating income, $109,000, as last year?

b. Assume the new plant is built and that next year the company manufactures and sells 32,000 balls (the same number as sold last year). Prepare a contribution format income statement and Compute the degree of operating leverage.

QUESTION

Whirly Corporation’s contribution format income statement for the most recent month is shown below:

Total and per unit

Sales (7,600 units) $258,400 $34.00

Variable expenses 144,400 19.00

Contribution margin 114,000 $15.00

Fixed expenses 55,300

Net operating income $58,700

Required: (Consider each case independently):

1. What would be the revised net operating income per month if the sales volume increases by 100 units?

2. What would be the revised net operating income per month if the sales volume decreases by 100 units?

3. What would be the revised net operating income per month if the sales volume is 6,600 units?

For each seperate "QUESTION" label it "ANSWER"
Answer #1

Feather Friends, Inc.

Sales units actual = 2880000/120 = 24000 units
Expected sales units = 24000 x 1.25 = 30000 units
Expected Sales price = $120 x 0.89 = $106.80

1. (a)

Sales $                3,204,000 =30000*106.8
Variable Expenses $                1,800,000 =30000*60
Contribution Margin $                1,404,000
Fixed Expenses $                   225,000 =160000+65000
Net Operating Income $                1,179,000

(b) Net operating income decreases by $101000 i.e. $1280000 - 1179000

2.

Sales $                3,600,000 =30000*120
Variable Expenses $                1,860,000 =30000*62
Contribution Margin $                1,740,000


Increase in advertisement can be = $1740000 - 160000 - 1280000 = $300,000

As per HOMEWORKLIB RULES, we are suppose to answer 1 question, i have answered 1, so kindly post others separately